OpenStudy (anonymous):

Find the area, if it exists, in the first quadrant above y=1, and bounded by the y-axis and the curve xy=1

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@thomaster @geerky42

geerky42 (geerky42):

Ok so you have \(x = \dfrac{1}{y}\): So from what I sketched, you have integral to take; \[\int_{1}^\infty\dfrac{dy}{y} ~~~=~~~ \lim_{a\to\infty}\int_1^a\dfrac{dy}{y}\]